# Enhancing In-Context Learning Performance with just SVD-Based Weight Pruning: A Theoretical Perspective

### v1

@@ 3.3 Generalization Bounds of ICL via Full Implicit GD Trajectories
It is notable that $\mathbf{G}_{t}$ is highly correlated with the sequence $(\Delta\mathbf{W}_{1},..,\Delta\mathbf{W}_{t})$. More precisely, adjusting the weight matrix of the $k$-th layer, will invariably impact the norm of $[\Delta\mathbf{W}_{t}]_{t>k}^{L}$, further influence $[\mathbf{G}_{t}]_{t>k}^{L}$. When we calibrate the weight matrix of the $k$-th layer, the span of affected weight updates $\Delta\mathbf{W}_{t}$ encompasses $L-k+1$ matrices, this indicates that the deeper the layer of the adjusted parameters is, the fewer the number of $\mathbf{G}_{t}$ affected. It also suggests that tweaks to the deep layers yield a more steadfast influence on the global norm of $[\mathbf{G}_{t}]_{1}^{L}$, thereby exerting a steadier influence over generalization performance. The enhanced stability observed in adjusting deeper layers, as demonstrated in our experimental findings in Section 2, can be theoretically explained based on the analysis presented above.
@@ 3.4 Applications of Our Theoretical Understanding
where $l_{1},l_{2}$ are coefficients and $\lambda_{1},\lambda_{2}$ are eigenvalues. Notice that if $\lambda_{1}\gg\lambda_{2}$, when point $\mathbf{b}$ moves in the direction of $\mathbf{x_{1}}$, the value of $l_{1}$ changes but the solution set $\mathbf{x}$ changes insignificantly. Conversely, if it moves in the direction of $\mathbf{x}_{2}$, the value of $l_{2}$ changes and the solution set $\mathbf{x}$ changes dramatically, this is an ill-conditioned444 Now consider a linear system $\mathbf{Wx}=\mathbf{b}$, where the solution set $\mathbf{x}$ is highly sensitive to the coefficients of $\mathbf{W}$ and $\mathbf{b}$. In such cases, the system of equations is termed ill-conditioned. problem whose solution is highly biased under small perturbations.
@@
Matrix condition number. $Cond(\mathbf{W})=||\mathbf{W}||_{p}||\mathbf{W}^{-1}||_{p}$. Specifically, the condition number $Cond(\mathbf{W})=\frac{\sigma_{max}}{\sigma_{min}}$ when $p=2$, where $\sigma_{max},\sigma_{min}$ denote the maximum and minimum singular values respectively. Generally, for matrices of the same order, the higher the condition number of a matrix is, the greater its ill-conditionedness will be.
@@
Based on all our exciting insights, we find it intuitive to design ICL improvements based on them, especially in downstream tasks. We propose a method where we first select layers with the top-k largest condition numbers and then identify the layer with the largest number among these. We perform a greedy search for the optimal clipping rate $\xi^{*}$ on the validation set and subsequently evaluate it on the test set. This procedure is reported in Algorithm 1 in Section C.2.
@@
Experimental results. As Figure 4 shows, the results indicate that Algorithm 1 can visibly influence performance across different language understanding tasks in ICL inference. Specially, the effectiveness of Algorithm 1 in enhancing performance varies not only by the model but also significantly by the task, indicating a potential task-specific and model-specific threshold for the benefits derived from algorithmic enhancements. More notably, Algorithm 1 is a derivative-free optimization method and compresses the model to some extent, providing developers and researchers with an effective approach for handling downstream tasks.
@@ 4 Conclusion and Limitation
In this paper, we show our surprising findings in ICL inference: SVD-based weight pruning can enhance ICL performance both in shallow and deep layers across different module types, and pruning weights in deep layers often results in more stable performance improvements in shallow layers. We conduct an in-depth theoretical analysis and explanation of these findings. Specifically, we first present the implicit gradient descent trajectories of ICL, afterwards, we give a mutual information based generalization bounds of ICL via full implicit GD trajectories. Based on all our experimental and theoretical insights, we intuitively propose a derivative-free and effective method for downstream tasks in enhancing ICL inference. However, further studies are required on (i) how to extend our generalization theory to a more standard Transformer architecture, (ii) do the results about ICL hold true for tasks beyond natural language processing and (iii) how to minimize the cost of searching the optimal clipping rate. Those will deepen our understanding of the ICL capabilities.
Appendix A Related Works
Low-rank properties of deep neural networks. In recent years, it has seen growing theoretical and experimental evidence of low-rank properties of deep neural networks. To name a few, Frankle and Carbin [14] indicate that neural networks can typically have over 90% of their weights eliminated with little to no loss in performance. Analogous results have been demonstrated in both feed-forward and convolutional networks used for image classification [22, 46, 20, 35]. More specifically, Sharma et al. [36] presents that careful pruning done at specific layers of Transformer models can produce significant boosts in performance on some tasks. The discovery of this phenomenon heightened interest in the connection between generalization and over-parameterization [51, 52], it also spurred research into developing pruning strategies that facilitate efficient model inference [28].
Appendix B Omitted Proofs and Additional Results
B.1 Proof of Lemma 1
@@
Given that numerous studies [19, 10, 33] describe ICL with attention can be connected to a meta-optimizer produces implicit gradient, we show the dual form between linear attention and gradient descent. First, consider a very simple linear layer,
@@
B.2 Softmax Attention and ICL Implicit GD
Inspired by Ren and Liu [33], by connecting Softmax Attention with Kernels, we can interpret ICL as a gradient descent process in contrast learning pattern. For simplicity, as in Section 3.1, scale factors are removed as follows:
@@ B.2 Softmax Attention and ICL Implicit GD
B.3 Feed-Forward and ICL Implicit GD
@@
where $\mathbf{W}_{MLP}\in\mathbb{R}^{(dout+din)\times(dim1)}$, which can be seen as a dimensional adaptation. And then do the similar derivation that we do in Appendix B.1, we can get $\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}=\mathbf{W}_{MLP}\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$.
B.4 Proof of Theorem 1
@@
Similar to B.1, Pay attention to the test token when $i=N+1$,
@@ B.5 Proof of Theorem 2
(Xu and Raginsky [50, Theorem 1.]). Assume the loss $\ell(w,\mathbf{h})$ is R-subGaussian for any $w\in\mathcal{W}$, then
@@
(Wang and Mao [47, Lemma 4.]). Let Assumption 1 hold, then $I(W_{L};\mathbf{H}_{S})\leq\sum_{t=1}^{L}I(-G_{t}+C_{t}^{1/2}N_{t};\mathbf{H}_{S}|W_{t-1})$. Let $-G_{t}+C_{t}^{1/2}N_{t}\triangleq\mathcal{G}_{t}$.
@@
(Polyanskiy and Wu [30, Corollary 3.1.]). For two random variables $X$ and $Y$, we have $I(X;Y)=\inf_{P}\mathbb{E}_{X}[D_{KL}(Q_{Y|X}||P)]$, where the infimum is achieved at $P=Q_{Y}$.
@@
(Wang and Mao [47, Lemma 5.]). At every time step $t$, let $P_{\tilde{N_{t}}|W_{t-1}}$be any distribution satisfying $D_{KL}(P_{\mathcal{G}_{t}|W_{t-1}}\|P_{\tilde{N_{t}}|W_{t-1}})<\infty$, we have $I(\mathcal{G}_{t};\mathbf{H}_{s}|W_{t-1})=\mathbb{E}_{W_{t-1}}\left[\inf_{P_{\tilde{N_{t}}|W_{t-1}}}\mathbb{E}_{\mathbf{H}_{S}}^{W_{t-1}}\left[D_{KL}(Q_{\mathcal{G}_{t}|\mathbf{H}_{s},W_{t-1}}\|P_{\tilde{N_{t}}|W_{t-1}})\right]\right]$, where the infimum is achieved when $P_{\tilde{N_{t}}|W_{t-1}}=Q_{\mathcal{G}_{t}|W_{t-1}}$. The KL divergence may be viewed as an estimate of the sensitivity of the full batch implicit gradient to a specific demonstration sample $\mathbf{H}_{s}=H_{s}$.
@@ B.6 Examples of Remark 4
$\left\lVert\mathbf{A+B}\right\rVert_{F}^{2}\leq\left\lVert\mathbf{A}\right\rVert_{F}^{2}+\left\lVert\mathbf{B}\right\rVert_{F}^{2}$ (20)
@@
Suppose we decompose $\mathbf{W}_{Q}$ by SVD, $\mathbf{W}_{Q}=\mathbf{W}_{Q_{r}}+\delta_{Q}=\mathbf{U}_{r}^{Q}\mathbf{\Sigma}_{r}^{Q}(\mathbf{V}_{r}^{Q})^{T}+\delta_{Q}$, let $\mathbf{z}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)$ for simplicity.
$\left\lVert\Delta\mathbf{W}_{icl}\right\rVert_{F}^{2}$ $=\left\lVert\mathbf{zU}_{r}^{Q}\mathbf{\Sigma}_{r}^{Q}(\mathbf{V}_{r}^{Q})^{T}+\mathbf{z}\delta_{Q}\right\rVert_{F}^{2}$
$\leq\left\lVert\mathbf{zW}_{Q_{r}}\right\rVert_{F}^{2}+\left\lVert\mathbf{z}\delta_{Q}\right\rVert_{F}^{2},$
@@
Suppose we decompose $\mathbf{W}_{V}$ by SVD, $\mathbf{W}_{V}=\mathbf{W}_{V_{r}}+\delta_{V}=\mathbf{U}_{r}^{V}\mathbf{\Sigma}_{r}^{V}(\mathbf{V}_{r}^{V})^{T}+\delta_{V}$, and consider the upper bound defined in Example 1,
@@
Suppose we decompose $\mathbf{W}_{MLP}$ by SVD, $\mathbf{W}_{MLP}=\mathbf{W}_{MLP_{r}}+\delta_{MLP}=\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{T}+\delta_{MLP}$, let $\mathbf{z}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$ for simplicity.
@@
$\left\lVert\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}\right\rVert_{F}^{2}$ $=\left\lVert\mathbf{U}_{r}\mathbf{\Sigma}_{r}\mathbf{V}_{r}^{T}\mathbf{z}+\delta_{MLP}\mathbf{z}\right\rVert_{F}^{2}$
$\leq\left\lVert\mathbf{W}_{MLP_{r}}\mathbf{z}\right\rVert_{F}^{2}+\left\lVert\delta_{MLP}\mathbf{z}\right\rVert_{F}^{2},$
@@
where recall Appendix B.3, $\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}=\mathbf{W}_{MLP}\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$. And the inequality part takes advantage of the Eq. (20) , so the upper bound on $\left\lVert\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}\right\rVert_{F}$ is decreasing when using SVD.

### v2

@@ 3.3 Generalization Bounds of ICL via Full Implicit GD Trajectories
It is notable that $\mathbf{G}_{t}$ is highly correlated with the sequence $(\Delta\mathbf{W}_{1},...,\Delta\mathbf{W}_{t})$. More precisely, adjusting the weight matrix of the $k$-th layer, will invariably impact the norm of $[\Delta\mathbf{W}_{t}]_{t>k}^{L}$, further influence $[\mathbf{G}_{t}]_{t>k}^{L}$. When we calibrate the weight matrix of the $k$-th layer, the span of affected weight updates $\Delta\mathbf{W}_{t}$ encompasses $L-k+1$ matrices, this indicates that the deeper the layer of the adjusted parameters is, the fewer the number of $\mathbf{G}_{t}$ affected. It also suggests that tweaks to the deep layers yield a more steadfast influence on the global norm of $[\mathbf{G}_{t}]_{1}^{L}$, thereby exerting a steadier influence over generalization performance. The enhanced stability observed in adjusting deeper layers, as demonstrated in our experimental findings in Section 2, can be theoretically explained based on the analysis presented above.
Remark 6 (How should Theorem 2 be interpreted?).
Expected generalization error (Theorem 2) = population risk ($L_{\mu}$) - empirical risk ($L_{\mathbf{H}_{s}}$). More specifically, on the one hand, Theorem 2 shows that clipping weights controls the F-norm of the implicit gradient $([\Delta\mathbf{W}_{t}]_{1}^{L}/[\mathbf{G}_{t}]_{1}^{L})$, which helps reduce the expected generalization error. On the other hand, we can evaluate the empirical risk $(L_{\mathbf{H}_{s}})$ by assessing the model’s performance on the validation set. If the generalization error is known, it is possible to estimate the population risk ($L_{\mu}$). Therefore, the most challenging aspect is addressing the generalization error. Furthermore, we will illustrate the interpretation of Theorem 2 from the perspectives of pruning methods and the ICL shot numbers. (i) Adjusting the F-norm of $([\Delta\mathbf{W}_{t}]_{1}^{L}/[\mathbf{G}_{t}]_{1}^{L})$ could enhance performance when utilizing ICL, suggesting that other weight-based pruning methods may also be effective. For example, magnitude-based pruning [52] directly controls the matrix F-norm. Certainly, there are also some layer-based pruning methods (e.g., drop-layer method [19]), we discuss in detail in Appendix B.6. (ii) Reducing the ICL shot numbers from $N$ to $N^{{}^{\prime}}$ can indeed lead to more robust outcomes specifically in the “SVD weight pruning” operation, rather than in the model’s overall performance. Experimentally, in Section 2.2, we analyze the effect of different ICL shot numbers. As shown in Figure 2, with a decrease in the number of shots, the rate of performance collapse in the model slows down after a sharp reduction at the shallow layer. Theoretically, $\Delta{\mathbf{W}}(N)-\Delta{\mathbf{W}}(N^{{}^{\prime}})=\left(\sum_{i=N^{{}^{\prime}}+1}^{N}{\mathbf{W}}_{V}{\mathbf{h}}_{i}\otimes{\mathbf{W}}_{K}\mathbf{h}_{i}\right)\mathbf{W}_{Q}$, suggesting that the implicit gradient for $N$ is expected to be more sensitive than that for $N^{{}^{\prime}}$.
@@ 3.4 Applications of Our Theoretical Understanding
where $l_{1},l_{2}$ are coefficients and $\lambda_{1},\lambda_{2}$ are eigenvalues. Notice that if $\lambda_{1}\gg\lambda_{2}$, when point $\mathbf{b}$ moves in the direction of $\mathbf{x_{1}}$, the value of $l_{1}$ changes but the solution set $\mathbf{x}$ changes insignificantly. Conversely, if it moves in the direction of $\mathbf{x}_{2}$, the value of $l_{2}$ changes and the solution set $\mathbf{x}$ changes dramatically, this is an ill-conditioned666 Now consider a linear system $\mathbf{Wx}=\mathbf{b}$, where the solution set $\mathbf{x}$ is highly sensitive to the coefficients of $\mathbf{W}$ and $\mathbf{b}$. In such cases, the system of equations is termed ill-conditioned. problem whose solution is highly biased under small perturbations.
@@
Matrix condition number [^7] Matrix condition number is an option, any indicator that can guide the control of norms is a potential option.. $Cond(\mathbf{W})=||\mathbf{W}||_{p}||\mathbf{W}^{-1}||_{p}$. Specifically, the condition number $Cond(\mathbf{W})=\frac{\sigma_{max}}{\sigma_{min}}$ when $p=2$, where $\sigma_{max},\sigma_{min}$ denote the maximum and minimum singular values respectively. Generally, for matrices of the same order, the higher the condition number of a matrix is, the greater its ill-conditionedness will be.
@@
Based on all our exciting insights, we find it intuitive to design ICL improvements based on them, especially in downstream tasks. We propose a method where we first select layers with the top-k largest condition numbers and then identify the layer with the largest number among these. We perform a greedy search for the optimal clipping rate $\xi^{*}$ on the validation set and subsequently evaluate it on the test set. This procedure is reported in Algorithm 1 in Appendix C.2.
@@
Experimental results. As Figure 4 shows, the results indicate that Algorithm 1 can visibly influence performance across different language understanding tasks in ICL inference. Specially, the effectiveness of Algorithm 1 in enhancing performance varies not only by the model but also significantly by the task, indicating a potential task-specific and model-specific threshold for the benefits derived from algorithmic enhancements. More notably, Algorithm 1 is a derivative-free optimization method and compresses the model to some extent, providing developers and researchers with an effective approach for handling downstream tasks. We also invite readers to refer to Appendices B.3, B.4 and B.5 for discussions on What would be the effect of pruning only a single module?, Why optimal clipping rate $\xi$ varies? and What would happen if we apply the same clipping rate to other datasets?
@@ 4 Conclusion and Limitation
In this paper, we show our surprising findings in ICL inference: SVD-based weight pruning can enhance ICL performance both in shallow and deep layers across different module types, and pruning weights in deep layers often results in more stable performance improvements than in shallow layers. We conduct an in-depth theoretical analysis and explanation of these findings. Specifically, we first present the implicit gradient descent trajectories of ICL, afterwards, we give a mutual information based generalization bounds of ICL via full implicit GD trajectories. Based on all our experimental and theoretical insights, we intuitively propose a derivative-free and effective method for downstream tasks in enhancing ICL inference. However, further studies are required on (i) how to extend our generalization theory to a more standard Transformer architecture, (ii) do the results about ICL hold true for tasks beyond natural language processing and (iii) how to minimize the cost of searching the optimal clipping rate. Those will deepen our understanding of the ICL capabilities.
Acknowledgments and Disclosure of Funding
This research was supported by National Natural Science Foundation of China (No.62476277, No.6207623), Beijing Natural Science Foundation (No.4222029), CCF-ALIMAMA TECH Kangaroo Fund(No.CCF-ALIMAMA OF 2024008), and Huawei-Renmin University joint program on Information Retrieval. We also acknowledge the support provided by the fund for building worldclass universities (disciplines) of Renmin University of China and by the funds from Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China, from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, from Public Policy and Decision-making Research Lab of Renmin University of China, and from Public Computing Cloud, Renmin University of China.
Appendix A Omitted Proofs and Additional Results
A.1 Proof of Lemma 1
@@
Given that numerous studies [21, 10, 35] describe ICL with attention can be connected to a meta-optimizer which produces implicit gradient, we show the dual form between linear attention and gradient descent. First, consider a very simple linear layer,
@@
A.2 Softmax Attention and ICL Implicit GD
Inspired by Ren and Liu [35], by connecting Softmax Attention with Kernels, we can interpret ICL as a gradient descent process in contrast learning pattern. For simplicity, as in Section 2, scale factors are removed as follows:
@@ A.2 Softmax Attention and ICL Implicit GD
A.3 Feed-Forward and ICL Implicit GD
@@
where $\mathbf{W}_{MLP}\in\mathbb{R}^{(dout+din)\times(dim1)}$, which can be seen as a dimensional adaptation. And then do the similar derivation that we do in Appendix A.1, we can get $\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}=\mathbf{W}_{MLP}\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$.
A.4 Proof of Theorem 1
@@
Similar to A.1, Pay attention to the test token when $i=N+1$,
@@ A.5 Proof of Theorem 2
(Xu and Raginsky [55, Theorem 1.]). Assume the loss $\ell(w,\mathbf{h})$ is R-subGaussian for any $w\in\mathcal{W}$, then
@@
(Wang and Mao [51, Lemma 4.]). Let Assumption 1 hold, then $I(W_{L};\mathbf{H}_{S})\leq\sum_{t=1}^{L}I(-G_{t}+C_{t}^{1/2}N_{t};\mathbf{H}_{S}|W_{t-1})$. Let $-G_{t}+C_{t}^{1/2}N_{t}\triangleq\mathcal{G}_{t}$.
@@
(Polyanskiy and Wu [32, Corollary 3.1.]). For two random variables $X$ and $Y$, we have $I(X;Y)=\inf_{P}\mathbb{E}_{X}[D_{KL}(Q_{Y|X}||P)]$, where the infimum is achieved at $P=Q_{Y}$.
@@
(Wang and Mao [51, Lemma 5.]). At every time step $t$, let $P_{\tilde{N_{t}}|W_{t-1}}$be any distribution satisfying $D_{KL}(P_{\mathcal{G}_{t}|W_{t-1}}\|P_{\tilde{N_{t}}|W_{t-1}})<\infty$, we have $I(\mathcal{G}_{t};\mathbf{H}_{s}|W_{t-1})=\mathbb{E}_{W_{t-1}}\left[\inf_{P_{\tilde{N_{t}}|W_{t-1}}}\mathbb{E}_{\mathbf{H}_{S}}^{W_{t-1}}\left[D_{KL}(Q_{\mathcal{G}_{t}|\mathbf{H}_{s},W_{t-1}}\|P_{\tilde{N_{t}}|W_{t-1}})\right]\right]$, where the infimum is achieved when $P_{\tilde{N_{t}}|W_{t-1}}=Q_{\mathcal{G}_{t}|W_{t-1}}$. The KL divergence may be viewed as an estimate of the sensitivity of the full batch implicit gradient to a specific demonstration sample $\mathbf{H}_{s}=H_{s}$.
@@ A.6 Examples of Remark 4
$\left\lVert\mathbf{A+B}\right\rVert_{F}\leq\left\lVert\mathbf{A}\right\rVert_{F}+\left\lVert\mathbf{B}\right\rVert_{F}$ (20)
@@
Suppose we decompose $\mathbf{W}_{Q}$ by SVD, $\mathbf{W}_{Q}=\mathbf{W}_{Q_{r}}+\delta_{Q}=\mathbf{U}_{:r}^{Q}\mathbf{\Sigma}_{:r}^{Q}(\mathbf{V}_{:r}^{Q})^{T}+\delta_{Q}$, let $\mathbf{z}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)$ for simplicity.
$\left\lVert\Delta\mathbf{W}_{icl}\right\rVert_{F}$ $=\left\lVert\mathbf{zU}_{:r}^{Q}\mathbf{\Sigma}_{:r}^{Q}(\mathbf{V}_{:r}^{Q})^{T}+\mathbf{z}\delta_{Q}\right\rVert_{F}$
$\leq\left\lVert\mathbf{zW}_{Q_{r}}\right\rVert_{F}+\left\lVert\mathbf{z}\delta_{Q}\right\rVert_{F},$
@@
Suppose we decompose $\mathbf{W}_{V}$ by SVD, $\mathbf{W}_{V}=\mathbf{W}_{V_{r}}+\delta_{V}=\mathbf{U}_{:r}^{V}\mathbf{\Sigma}_{:r}^{V}(\mathbf{V}_{:r}^{V})^{T}+\delta_{V}$, and consider the upper bound defined in Example 1,
@@
Suppose we decompose $\mathbf{W}_{MLP}$ by SVD, $\mathbf{W}_{MLP}=\mathbf{W}_{MLP_{r}}+\delta_{MLP}=\mathbf{U}_{:r}\mathbf{\Sigma}_{:r}\mathbf{V}_{:r}^{T}+\delta_{MLP}$, let $\mathbf{z}\triangleq\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$ for simplicity.
@@
$\left\lVert\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}\right\rVert_{F}$ $=\left\lVert\mathbf{U}_{:r}\mathbf{\Sigma}_{:r}\mathbf{V}_{:r}^{T}\mathbf{z}+\delta_{MLP}\mathbf{z}\right\rVert_{F}$
$\leq\left\lVert\mathbf{W}_{MLP_{r}}\mathbf{z}\right\rVert_{F}+\left\lVert\delta_{MLP}\mathbf{z}\right\rVert_{F},$
@@
where recall Appendix A.3, $\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}=\mathbf{W}_{MLP}\left(\sum_{i=1}^{N}\mathbf{W}_{V}{\mathbf{h}_{i}}\otimes\mathbf{W}_{K}{\mathbf{h}_{i}}\right)\mathbf{W}_{Q}$. And the inequality part takes advantage of the Eq. (20) , so the upper bound on $\left\lVert\Delta\mathbf{W}_{icl}^{{}^{\prime\prime}}\right\rVert_{F}$ is decreasing when using SVD.
Appendix B More Discussions
B.1 Explanation of the Mask Matrix
Our notation here follows the related work [2], which explains: Note that the prompt is asymmetric since the label for $\mathbf{x}_{N+1}$ is excluded from the input. To reflect this asymmetric structure, the mask matrix $\mathbf{M}$ is included in the attention. More specifically, if you pay attention to the $(N+1)$-th item, $\mathbf{M}$ is supposed to represent a causal mask (For $\mathbf{H}\in\mathbb{R}^{(dout+din)\times(N+1)}$, $\mathbf{HM}=(\mathbf{h}_{1},...,\mathbf{h}_{N},\mathbf{h}_{N+1})\mathbf{M}=(\mathbf{h}_{1},...,\mathbf{h}_{N},0)=\mathbf{H}_{s}$). Besides, this mask method is used in GLM training [13]: Part A tokens can attend to each other, but cannot attend to any tokens in B. Part B tokens can attend to Part A and antecedents in B, but cannot attend to any subsequent tokens in B. So it is reasonable for Eq.(1) and Eq.(2) use the same mask.
B.2 Definitions of Deep and Shallow
Indeed, there is no universally accepted definition for the terms “deep” and “shallow” as their interpretation can be subjective and dependent on the reference frame (e.g., model size). Intuitively, in this paper, our definition is similar to that of work [50]: “shallow” layers refer to those closer to the input, while “deep” layers are closer to the output. In Figure 1 and Figure 2 (Section 2), "shallow" typically denotes the first few layers, and "deep" denotes the last few layers of the network.
B.3 The Effect of Pruning Only a Single Module
In our theory, the example (pruning only $\mathbf{W}_{K}$ or $\mathbf{W}_{V}$) is provided in Appendix A.6, demonstrating how weight pruning can affect the norm of $[\Delta\mathbf{W}_{t}]_{1}^{L}$/$[\mathbf{G}_{t}]_{t}^{L}$. Thus, it may confer advantages on the performance of Transformers in ICL inference. However, in Theorem 2 (Remark 5), it also suggests that modifications to the shallow layers have a less steady impact. Additionally, please review the supplementary experimental results provided below. (We choose key projection matrix $\mathbf{W}_{K}$ and select the 3-layer (large matrix condition number in Figure 3) of GPT-J-6B, others are the same to Section 2.1).
Task Name Optimal $\xi^{*}$ Test Accuracy/F1 Improve
SST-2 0.0 0.7828 - 0.7828
RTE 0.5 0.5413 - 0.5413
COPA 0.995 0.53 - 0.54
B.4 Why Optimal Clipping Rate Varies?
As we mentioned in Section 3.3 and Section C.4, the implicit gradients produced by Transformers in practical applications are noisy due to factors such as the extent of model pre-training and data characteristics (e.g., ICL shot number/task difficulty). Therefore, $[\Delta\mathbf{W}_{t}]_{1}^{L}$/$[\mathbf{G}_{t}]_{t}^{L}$ in Theorem 1 have varies noise. That is why optimal $\xi$ varies.
Gradient quality derived from context (i.e., $\mathbf{G}_{t}$).
  • In Theorem 1 (Remark 2), $\mathbf{G}_{t}$ is only dependent on $\mathbf{W}_{t-1}$ and $\mathbf{H}_{s}$, this is consistent with gradient descent in terms of relevance (Conventionally, gradients in training are only related to the current parameters and the training samples).
  • In real-world training scenarios, SGD computes gradient by selecting a small batch of samples per iteration. This approach approximates the true gradient, inherently introducing noise. Similarly, in In-Context Learning, a small subset of samples (context examples) is used to generate implicit gradient (i.e., $\mathbf{G}_{t}$), which also results in the introduction of noise.
B.5 Apply the Same Clipping Rate to Other Datasets
As we mentioned in Appendix B.4, $[\Delta\mathbf{W}_{t}]_{1}^{L}$/$[\mathbf{G}_{t}]_{t}^{L}$ in Theorem 1 exhibit varying levels of noise, causing the optimal clipping rate to vary among different tasks, as it is dependent on the specific task and data. So if we apply the clipping rate (0.95) as used in SST-2 to other datasets, the model performance can either improve or deteriorate. Additionally, It is possible to conduct a certain number of experiments to find a range of optimal clipping rate that is broadly applicable.
B.6 What Would Happen if the Layer Was Dropped Entirely?
Dropping the layer could be the best option specifically for optimizing generalization error. Here’s a detailed analysis:
(1) In our theoretical framework, we model each layer of the Transformer do a single iteration of implicit gradient descent (ICL) in Theorem 1. This scientific analysis references [46, 3, 10].
(2) In Theorem 2, L-layer Transformer:
Expected generalization error = population risk ($L_{\mu}$) - empirical risk ($L_{\mathbf{H}_{s}}$)
Expected generalization error = $\sqrt{\frac{R^{2}}{N}\sum_{t=1}^{L}d\log(\left\lVert\Delta\mathbf{W}_{t}\right\rVert_{F}^{2}\left\lVert\mathbf{G}_{t}\right\rVert_{F}^{2})}$ (only show the main part). If you drop the entire layer, it will change from $\sum_{t=1}^{L}$ to $\sum_{t=1}^{L-1}$. Therefore, dropping the layer may in fact be the best option for generalization error.
(3) However, according to the traditional statistical learning viewpoint, performance can be bounded by the sum of optimization error and generalization error (Please see Remark 6 for “How should Theorem 2 be interpreted?”).
Thus, during the pruning process, there is a trade-off between optimization and generalization. Therefore, as pruning increases, the model’s performance tends to first improve and then decline (the drop-layer method [19] does not harm but also does not improve), as demonstrated in the experiments shown in Figure 1.
In conclusion, dropping the entire layer can be a potential method (best option for generalization error) in our theoretical framework, but it may not necessarily be the best option for model performance.